OpenStudy (anonymous):

The population of a certain city was 123,000 in 2006, and the observed relative growth rate is 2% per year. (a) Find a function that models the population after t years. (Assume t = 0 represents the year 2006.) (b) Find the projected population in the year 2012. (Round your answer to the nearest thousand.) (c) In what year will the population reach 200,000?

OpenStudy (anonymous):

\[123,000\times (1.02)^t\]

OpenStudy (anonymous):

\[123,000\times (1.02)^6=138518\]

OpenStudy (anonymous):

Congrats , 3000 medals

OpenStudy (anonymous):

wow really? i remember the good old days when i was counting down amistre to 1000!

OpenStudy (anonymous):

\[200,000=123,000(1.02)^t\] \[\frac{200}{123}=(1.02)^t\] \[t=\frac{\ln(\frac{200}{123})}{\ln(1.02)}\]

OpenStudy (anonymous):

about 24.5 years

OpenStudy (anonymous):

so 2031?